# DivScene: Benchmarking LVLMs for Object Navigation with Diverse Scenes and Objects

### v2

@@ 4.3 Statistics of Both Datasets
In DivScene, we collected 4,614 houses across 81 scene types. To the best of our knowledge, this dataset covers the widest range of scene types. We show the diversity of our collected houses in Figure 7 by plotting most types under each category. Thanks to Objaverse assets (Deitke et al., 2023), the collected houses contain objects from 22,696 different types, including very common objects such as fridges, beds, shelves, and sofas, and rare objects such as multicolored bookshelf and vintage wooden bench.
@@
Then, we randomly pick one house from each scene type to build a test set of 81 houses. Similarly, we randomly select 27 houses from distinct scene types as the validation set. Thus, DivScene is divided into training, validation, and test sets, covering 4506, 27, and 81 houses. On the training set, we sample five episodes in each scene with different target objects. Four episodes are selected in each house in evaluation sets to balance the evaluation efficiency and accuracy. In total, the DivTraj dataset contains 22962 episodes and 5707 different kinds of target objects. More statistics of the DivScene and DivTraj are shown in Section A.6.
@@ 5 NatVLM
Figure 1 illustrates the datils of the proposed NatVLM model. Here, we build an end-to-end agent that imitates shortest path experts in AI2THOR. Our agent is tuned from the LVLM, Idefics 2 (Laurençon et al., 2024), in an end-to-end manner. It takes an environment observation at each time step $t$ and generates the next action $a_{t}$, following the same format as instruction tuning. Here, we discuss the collection of instructions and corresponding responses.
5.1 Instructions Compilation
@@
We manually build instructions for each time step in an episode, which contains four parts. First, the instruction provides a brief introduction to the object navigation task, such as possible actions and step length. Then, we add the episode-specific information, including the target object, its positions, the agent’s position and rotation, and visual observation. Third, we provide the agent’s positions and actions for the recent $M=8$ steps, along with the visual observations from the recent $K=4$ steps, to help the agent make better decisions. Finally, the instruction asks the model to predict the next action by considering the agent’s position and visual observation in accordance with the CoT explanation in the responses. We show the specific instruction template in Section B.1.
5.2 Response Collection with CoT Explanation Traces
@@
We tune the LVLM to generate the next action when instructions are given. In the model, we encode actions as natural language and use the model’s generative capabilities to decode the next action. Nonetheless, we found that merely requiring LVLMs to output the next action leads to unsatisfactory results. The model only grasps the surface-level styles of the prompt but misses the underlying rationales of object navigation.
@@
To enhance the agent’s understanding of navigation rationale, we manually build responses with CoT explanation traces during the navigation process. The structure of the responses covers three steps. In the first step, we have the agent compare its current position with the target and determine whether it needs to move forward or take another action. After this, the agent is asked to check the obstacles in the visual observation to see whether it needs to rotate. In the last step, the agent gives the final decision based on the analyses in the first two steps. In contrast to writing explanations with LLMs (Mitra et al., 2023), we manually write the prompt template of explanation traces for the MoveAhead and Done actions, leaving the position information to be filled in with coordinates at each step. For the rotation actions (i.e., RotateRight and RotateLeft), we identify very common scenarios during the navigation and design the explanation trace template for each of them. Then, we introduce a few postprocessing steps, like action balancing and conflict filtering. We show concrete prompts and postprocessing details in Sections B.2 and B.3, respectively.
@@ 6 Experiment
In this section, we conduct extensive experiments to compare our framework with various baselines.
6.1 Dataset and Metrics
We conduct our experiments on DivTraj with the statistics shown in Section 4.3. Our agent NatVLM and baselines navigate in a given house until the model chooses the Done action or reaches the max action limit, 200 in our experiments. We report the metrics Success Rate (SR), Success weighted by Path Length (SPL (Anderson et al., 2018)), and Success weighted by Episode Length (SEL (Eftekhar et al., 2023)). An episode is considered successful when the target object appears in the agent egocentric observation and is less than 1.5 meters away. Specifically, SR and SPL are computed as $\frac{1}{N}\sum_{i=1}^{N}S_{i}$ and $\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{max(l_{i},p_{i})}$, where $N$ is the number of episodes, $S_{i}$ is the indicator of success, $l_{i}$ is the length of the shortest path, and $p_{i}$ is the length of the predicted trajectory. For SEL, we replace $l_{i}$ and $p_{i}$ with the action number of the shortest and predicted paths.
Table 1: Performance of NatVLM and baselines on the validation and test sets of DivTraj. We also provide the performance on all evaluation data. The highest scores are bolded. In the “LLMs w/Captions” baseline, we use Llava 1.5 as the captioning model.
@@ 6.1 Dataset and Metrics
6.2 Baseline Methods
We compare our embodied agent with four baselines:
Blind LLMs: The first baseline we test is the text-only LLM agent that simply predicts the next action based on the textual instruction without considering any visual information. This agent references how far we can get solely using prior world knowledge and random guessing. For the LLM choice, we evaluate Llama 2 (7B, 13B) (Touvron et al., 2023), Llama 3.1 (8B) (Dubey et al., 2024), and Mistral (7B) (Jiang et al., 2023).
Socratic LLMs w/ Image Captions: This is the simplest agent that leverages perceptual information. Here, we use an image captioning model to convert visual observations into natural language. These captions provide a language description of egocentric images, allowing LLMs to obtain the content of perceptual information. Here, we employ Llava 1.5 (Liu et al., 2024a) as the captioning model while using the same LLMs as those in the “Blind LLM” baseline.
Open-Source LVLMs: The most generic agent for navigation is one that can directly process textual instructions and visual observations. Thus, we directly test recent open-source LVLMs without any further tuning, which are capable of processing images in addition to textual queries. Here, the single image LVLM we test is Llava 1.5 (7B, 13B) (Liu et al., 2024b; a). Meanwhile, we test Qwen-VL (7B) (Bai et al., 2023) and Idefics 2 (8B) (Laurençon et al., 2024) for handling multiple images.
API-based LVLMs: In addition to open-source LVLMs, we also evaluate closed-source ones, including GPT-4v (OpenAI, 2023) and GPT-4o (OpenAI, 2024). They can process multiple images and achieve state-of-the-art performance in multimodal tasks.
6.3 Main Evaluation
We present the results of our model NatVLM and baselines on the validation and test sets in Table 1. We also include the performance of selecting a random action at each step (i.e., Random) as a reference. In general, our embodied agent NatVLM achieves the best performance on object navigation, exceeding the performance of all baselines by a large margin. For example, our model can successfully navigate to 57.41% of episodes on the test set, increasing by about 20% compared to the GPT-4o baseline. Meanwhile, according to the higher SPL and SEL scores on both test and validation sets, our model can navigate to target objects with better efficiency.
For the baselines, the blind LLMs achieve performance slightly higher than the random results. For example, LLama 3.1 (8B) achieves a success rate of 11.57%, about 4 points higher than the random guess. In the meantime, we observe that the performance of all LLMs only improved marginally when we added the captioning model to provide additional perceptual information. This result shows that the captioning model can miss important image content details, leading to unsatisfactory improvement. On the other hand, we find that LVLMs can achieve the best results across all baselines. For example, the closed-source VLMs, GPT-4v and GPT-4o, can successfully navigate to target objects in more than 30% cases. In addition, Idefics 2 (8B) attains success rates exceeding 20% on both validation and test sets.
Table 2: The ablation study of our agent NatVLM. $\Delta_{*}$ columns indicate the score difference of each metric on the test set. We remove the explanation trace and test different methods of position comparisons. We bold the highest scores except for the gold label test ($\diamond$ w Gold Label).
@@ 6.3 Main Evaluation
(a) Performance of the default prompt.
(b) Performance of the difference-equation prompt
Figure 4: Design investigation. We provide different numbers of recent visual observations with the embodied agent. Besides the default prompt we use, we also evaluate the difference-equation prompt. See scores of all metrics in Section C.2
6.4 Ablation Study
@@
To better understand the role of CoT explanation traces in our agent NatVLM, we conduct two ablation studies to analyze its contribution. First, we verify the efficacy of CoT explanation traces. Then, we analyze different ways to compare the positions of the agent and target object. The results of experiments are shown in Table 2.
@@
First, we remove the whole explanation traces in the response of the instruction tuning data ($\diamond$ w/o ET). Thus, we fine-tune the agent to only generate the next action. From the result, we can find that the performance of our agent drastically drops. This verifies that explanation traces can help the LVLM to better understand the underlying rationales of object navigation. Subsequently, we further enhance the agent’s input by providing the gold label indicating the positional difference between the agent and the target object ($\diamond$ w/o ET & w Gold). As shown in Table 2, we can observe that the gold labels cannot help much as the performance only fluctuates somewhat.
@@
Then, we study the effects of the position comparison, a crucial component and the initial step in our explanation traces. The default prompt is to directly generate the position difference: “the difference to the target object is [position_diff]” as shown in Table 10, where [position_diff] is a placeholder. First, we provide the global label of positional differences in the input instructions ($\diamond$ w Gold Label). The results in Table 2 show that providing global labels can improve the performance of our agent. This observation suggests that our agent may occasionally compute the positional difference with inaccuracies, as providing the gold label can lead to a performance increase. Next, we test the difference-equation prompt ($\diamond$ w Diff-EQ), where we fine-tune our agent to generate an equation for computing the positional difference. However, writing the equation of computation only leads to small variations in performance. All the abovementioned prompts are shown in Section C.1.
@@ 6.5 Design Investigation
In this experiment, we thoroughly investigate a few design decisions regarding the image number and step number of recent positions and actions. Besides the default prompt structure, we also conduct experiments with the difference-equation prompt, which is introduced as $\diamond$ w Diff-EQ” in the ablation study.
@@
The default design provides NatVLM with four images. In addition, we test the agent’s performance using different numbers of input images, including 2, 6, and 8 images. The results are plotted in Figure 4. First, we observe a decline in the agent’s performance when provided with only two images, showing that using fewer images leads to worse performance. For instance, our agent finishes the navigation successfully only 50% of the time, underperforming the 4-image baseline. Moreover, increasing the number of images to 6 or 8 does not result in further performance variations. Thus, we choose to provide our agent with four images for the tradeoff between accuracy and efficiency.
@@
We also investigate the effect of recent positions and actions on navigation performance. By default, we provide the agent with information about the recent 8 steps. Then, we test the performance when we provided the positions and actions of 4, 12, and 16 steps. As illustrated in Table 3, a substantial performance improvement is evident when increasing the number of steps from 4 to 8. However, further increases in step count yield limited returns. Thus, we provide our agent with 8 steps.
@@
Table 3: Hyperparameter investigation. Our agents receive various numbers of recent actions and positions within the default prompt. See results of difference-equation prompt in Section C.3.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
4 steps 46.30 38.66 39.82 45.99 37.45 39.60 46.14 38.05 39.71
8 steps 57.41 47.84 47.90 54.94 44.45 45.83 56.17 46.15 46.86
12 steps 42.59 35.92 36.43 50.93 41.15 43.20 46.76 38.53 39.81
16 steps 51.85 43.01 42.72 53.40 43.03 44.08 52.62 43.02 43.40
Table 4: Few-shot learning ability. We test our agent with different proportions of training data. Besides the default prompt, we also evaluate the difference-equation prompt in Section C.4.
% Data Valid Test Test Diff w GPT-4o
SR SPL SEL SR SPL SEL $\Delta_{SR}$ $\Delta_{SPL}$ $\Delta_{SEL}$
6.6 Few-shot Learning Ability
@@
Our agent undergoes instruction tuning with an extensive training dataset, DivTraj. To assess its generalization capabilities, we design a few-shot experiment to confirm its ability to generalize with fewer data. While the original training data contains five episodes per house, we train the model with only 1, 2, 3, and 4 episodes, representing 20%, 40%, 60%, and 80% of the full data. The results are shown in Table 4. For clarity, we also compare our models with the GPT-4o baseline. We can find that our agent has strong few-shot learning abilities. With only 20% percent training data, our agent can perform similarly to GPT-4o and generalize well on unseen houses. Meanwhile, the results demonstrate a gradual performance improvement as we incrementally increase the data volume, with performance gains plateauing at approximately 80% of the full dataset.
Table 5: Performance of zero-shot transfer learning on iTHOR and ProcTHOR. NatVLM outperforms all baselines by a large margin.
@@
NatVLM $\diamond$ w Diff-EQ 72.32 58.98 62.35 54.59 45.53 46.80
6.7 Zero-shot Transfer Learning
To further verify the generalization ability of NatVLM, we test it on other house datasets with limited scene types: iTHOR (Weihs et al., 2021) and ProcTHOR (Deitke et al., 2022). Both datasets encompass four distinct scene types: bedrooms, living rooms, kitchens, and bathrooms. There are 30 rooms in iTHOR for each scene type, which is designed by professional 3D artists. Meantime, ProcTHOR is a procedural house-generation system that constructs 10,000 unique houses automatically. Similar to iTHOR, we sample 30 houses for each scene type from ProcTHOR. Four episodes are sampled in each house to evaluate our agent.
Figure 5: Error analysis. The agent needs to find a soda can in the lower right corner of a game room.
We directly use NatVLM tuned on the DivTraj dataset to test this zero-shot transferring ability. Since we do not tune hyper-parameters on these datasets, we treat each dataset as a test set without any validation set. The results are shown in Table 5. We also report the performance of open-source VLMs and some ablated models as baselines. The results show that our agent surpasses all the baselines on both datasets. Such improvements indicate that our framework has a strong ability to generalize in other environments.

### v3

@@ 4.3 Statistics and Comparison
In DivScene, we collected 4,614 houses across 81 scene types. To the best of our knowledge, this dataset covers the widest range of scenes. We show the diversity of our collected houses in Figure˜7 by plotting most types under each category. Thanks to Objaverse (Deitke et al., 2023b), the collected houses contain objects from 22,696 different types, including very common objects such as fridges, beds, shelves, and sofas, and rare objects such as multicolored bookshelf and vintage wooden bench.
@@
Then, we randomly pick one house of each scene type to build a test set of 81 houses. Similarly, we randomly select 27 houses of distinct scene types as the validation set. DivScene is divided into training, validation, and test sets, covering 4,506, 27, and 81 houses. On the training set, we sample five episodes in each scene with different target objects. Four episodes are selected in each house in the evaluation sets to balance the evaluation efficiency and accuracy. In total, the DivSceneep dataset contains 22,962 episodes and 5,707 different kinds of target objects.
We provide a detailed comparison of scene complexity between our dataset and the other two datasets: iTHOR Weihs et al. (2021) and ProcTHOR Deitke et al. (2022). As shown in Table˜1, our dataset includes the most objects and scenes, and room sizes and numbers in our dataset remain consistent with manually created ones in iTHOR. More statistics are shown in Section˜A.6.
@@ 5 NatVLM
In this section, we describe our NatVLM model. It is fine-tuned from Idefics 2 (Laurençon et al., 2024) through imitation learning on shortest paths from DivSceneep to enhance the navigation ability. At each time step $t$, the model processes environment observations $s_{t}$ and generates the next action $a_{t}$ in a manner consistent with instruction tuning. Figure˜1 provides an overview of NatVLM.
5.1 Instruction Compilation
@@
We manually build instructions for each time step in an episode, which contains four parts: (i) the instruction provides a brief introduction to the object navigation task, such as possible actions and step length; (ii) we add the current environment observation, including the target object, its position, the agent’s position and rotation, and visual observation; (iii) we provide the agent’s positions and actions for the recent $M=8$ steps, along with the visual observations from the recent $K=4$ steps, to help the agent capture navigation history; (iv) the instruction asks the model to predict the next action by considering the current observation and navigation history in accordance with the CoT explanation in the responses. We show the specific instruction template in Section˜B.1.
5.2 Response with CoT Explanation
@@
We tune the Idefics 2 to generate the next action when instructions are given. Specifically, we encode actions as natural language and use the model’s generative capabilities to decode the next action. Nonetheless, we found that merely requiring LVLMs to output the next action leads to unsatisfactory results. The model only grasps the surface-level styles of the prompt but misses the underlying rationales of object navigation.
@@
To enhance its understanding of the object navigation task, we manually build responses with CoT explanation traces during the navigation process. The structure of the responses covers three steps. In the first step, we have the agent compare its current position with the target and determine whether it needs to move forward or take other actions. After this, the agent is asked to check the obstacles in the visual observation to see whether it needs to rotate. In the last step, the agent gives the final decision based on the analyses in the first two steps. In contrast to writing explanations with LLMs (Mitra et al., 2023), we manually write the prompt template of explanation traces, leaving the position information to be filled in with coordinates at each step. Meanwhile, we employ a few postprocessing steps, like action balancing and conflict filtering. We show concrete prompts and postprocessing details in Sections˜B.2 and B.3, respectively.
5.3 Imitation Learning Objective
With collected instructions and responses, we employ imitation learning to train NatVLM to mimic decision-making from shortest-path demonstrations. Specifically, we adopt the Behavior Cloning (BC) framework, where the goal is to learn a policy $\pi_{\theta}(a_{t}|s_{t})$ that maps the current state $s_{t}$ (environment observation) to the action $a_{t}$ at each time step $t$. We minimize the negative log-likelihood of the shortest-path actions over collected observation-action pairs $(s_{t},a_{t})$ obtained from trajectories in DivSceneep, which is denoted as $\mathcal{D}$. The objective function is given by:
$\mathcal{L}_{\text{BC}}(\theta)=-\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(a_{t}|s_{t})\big{]},$ (1)
where $\theta$ represents the learnable parameters of Idefics 2. As we aforementioned, the state $s_{t}$ is included in the instruction, and the action $a_{t}$ includes the CoT explanation besides the next action.
@@ 6 Experiment
We conduct extensive experiments to evaluate both off-the-shelf LVLMs and LLMs, as well as models that have been fine-tuned using DivSceneep.
6.1 Metrics
A model navigates in a house until it chooses the Done action or reaches the maximum action limit of 200 steps. We report the metrics Success Rate (SR), Success weighted by Path Length (SPL; Anderson et al., 2018), and Success weighted by Episode Length (SEL; Eftekhar et al., 2023). An episode is considered successful when the target object appears in the agent egocentric observation and is less than 1.5 meters away. Specifically, SR and SPL are computed as $\frac{1}{N}\sum_{i=1}^{N}S_{i}$ and $\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{max(l_{i},p_{i})}$, where $N$ is the number of episodes, $S_{i}$ is the indicator of success, $l_{i}$ is the length of the shortest path, and $p_{i}$ is the length of the predicted trajectory. For SEL, we replace $l_{i}$ and $p_{i}$ with the action number of the shortest and predicted paths.
@@ 6.1 Metrics
Table 2: Performance of all models DivSceneep. “All” means the average of both evaluation sets. The highest scores are bolded. In the “LLMs w/Captions” baseline, we use Llava 1.5 as the captioning model.
6.2 Evaluated Methods
Besides NatVLM, we comprehensively evaluate four off-the-shelf methods and one fine-tuned method on our sampled dataset [^2] We do not include complicated frameworks built on LVLMs as they are out of the evaluation scope. More in Section C.5.:
Blind LLMs are text-only LLMs that simply predict the next action based on the textual instruction without considering any visual information. This method references how far we can get solely using prior world knowledge and random guessing. For the LLM choice, we evaluate Llama 2 (7B, 13B) (Touvron et al., 2023), Llama 3.1 (8B) (Dubey et al., 2024), and Mistral (7B) (Jiang et al., 2023).
Socratic LLMs w/ Image Captions is the simplest method that leverages visual information. Here, we use an image captioning model to convert egocentric images into language descriptions, allowing LLMs to obtain the content of visual information. We employ Llava 1.5 (Liu et al., 2024a) as the captioning model while using the same LLMs as those in the “Blind LLM” baseline.
Open-Source LVLMs are directly tested without any further tuning. They are capable of processing images in addition to textual queries. Here, we test Llava 1.5 (7B, 13B) (Liu et al., 2024b, a), Qwen-VL (7B) (Bai et al., 2023), Idefics 2 (8B) (Laurençon et al., 2024), Qwen2.5-VL (7B) (Bai et al., 2025), InternVL3 (8B) (Zhu et al., 2025), and Gemma3 (12B) (Team et al., 2025).
API-based LVLMs we evaluated include GPT-4v (OpenAI, 2023) and GPT-4o (OpenAI, 2024). They can process multiple images and achieve state-of-the-art performance on multimodal tasks.
Fine-Tuned LVLM: For fine-tuned LVLMs on our data, we also evaluated fine-tuning Idefics 2 (Laurençon et al., 2024) on our sampled trajectories without any CoT explanation traces.
@@ 6.2 Evaluated Methods
Table 3: The ablation study. $\Delta_{*}$ columns show score differences. We remove explanation traces and test different methods of position comparisons. We bold the highest scores except for the gold label test ($\diamond$ w Gold Label).
6.3 Main Evaluation
We present the results of all off-the-shelf and fine-tuned methods on the validation and test sets in Table˜2. We also include the performance of selecting a random action at each step (i.e., Random) as a reference. In general, the fine-tuned NatVLM achieves the best performance on object navigation, exceeding other methods by a large margin. For example, NatVLM can successfully navigate to 57.41% of episodes on the validation set, increasing by about 20% compared to the GPT-4o baseline. Meanwhile, according to the higher SPL and SEL on both test and validation sets, NatVLM can navigate to target objects with better efficiency.
For the off-the-shelf methods, the blind LLMs achieve performance slightly higher than the random results. For example, Llama 3.1 (8B) achieves a success rate of 11.57%, about 4 points higher than the random guess. In the meantime, we observe that the performance of all LLMs only improved marginally when we added the captioning model to provide additional perceptual information. This result shows that the captioning model can miss important image content details, leading to unsatisfactory improvement. On the other hand, we find that LVLMs can achieve the best results across all baselines. For example, the closed-source LVLMs, GPT-4v and GPT-4o, can successfully navigate to target objects in more than 30% cases. In addition, Idefics 2 (8B) attains success rates exceeding 20% on both validation and test sets.
6.4 Ablation Evaluation
@@
To better understand the role of CoT explanation traces in NatVLM, we conduct two ablation studies to analyze its contribution. First, we verify the efficacy of CoT explanation traces. Then, we analyze different ways to compare the positions of the agent and target object. The results of experiments are shown in Table˜3.
@@
First, we remove explanation traces in the instruction tuning data ($\diamond$ w/o ET). Thus, we fine-tune the agent to only generate the next action. We find that the performance of our agent drastically drops, verifying that explanation traces can help LVLMs to better understand the underlying rationales of object navigation. Furthermore, we enhance the agent’s input by providing the gold label of the positional difference between the agent and the target object ($\diamond$ w/o ET & w Gold). We observe that the gold labels cannot help much, as the performance only fluctuates somewhat.
@@
Then, we study the effects of the position comparison, the crucial component in our explanation traces. The default prompt is to directly generate the position difference: “the difference to the target object is [position_diff]” as shown in Table˜11, where [position_diff] is a placeholder. First, we provide the global label of positional differences in the input instructions ($\diamond$ w Gold Label). The results in Table˜3 show that providing global labels can improve the performance of our agent, suggesting that our agent may occasionally compute the positional difference with errors. Next, we test the difference-equation prompt ($\diamond$ w Diff-EQ), where we fine-tune our agent to generate an equation for computing the positional difference. However, writing the equation of computation only leads to small variations in performance. All the abovementioned prompts are shown in Section˜C.1.
Figure 4: Design investigation. We provide different numbers of recent visual observations with the embodied agent. See scores of all metrics in Section˜C.2
@@ 6.5 Design Investigation
In this experiment, we thoroughly investigate a few design decisions regarding the image number and step number of recent positions and actions. We also test NatVLM with different prompts shown in Sections˜C.2 and C.3
@@
The default design provides NatVLM with four images. In addition, we test its performance using different numbers of input images, including 2, 6, and 8 images. The results are plotted in Figure˜4. First, we observe a decline in the agent’s performance when provided with only two images, showing that using fewer images leads to worse performance. For instance, our agent finishes the navigation successfully only 50% of the time, underperforming the 4-image baseline. Moreover, increasing the number of images to 6 or 8 does not result in further improvements. Thus, we choose to provide NatVLM with four images for the tradeoff between accuracy and efficiency.
@@
We also investigate the effect of recent positions and actions on navigation performance. By default, we provide NatVLM with information about the recent 8 steps. Here, we test the performance when we provide the positions and actions of 4, 12, and 16 steps. As illustrated in Table˜4, a substantial performance improvement is evident when increasing the number of steps from 4 to 8. However, further increases in step count yield limited returns. Thus, we provide our agent with 8 steps.
Number Valid Test
SR SPL SEL SR SPL SEL
4 steps 46.30 38.66 39.82 45.99 37.45 39.60
8 steps 57.41 47.84 47.90 54.94 44.45 45.83
12 steps 42.59 35.92 36.43 50.93 41.15 43.20
16 steps 51.85 43.01 42.72 53.40 43.03 44.08
@@
Table 4: Our agents receive various numbers of recent actions and positions. See all metrics in Section˜C.3.
% Data Test Test Diff w GPT-4o
SR SPL SEL $\Delta_{SR}$ $\Delta_{SPL}$ $\Delta_{SEL}$
20 38.89 31.40 32.13 $\uparrow$0.62 $\downarrow$0.34 $\uparrow$4.21
40 38.27 31.00 30.79 $\downarrow$0.00 $\downarrow$0.74 $\uparrow$2.87
60 52.12 42.25 44.40 $\uparrow$13.85 $\uparrow$10.51 $\uparrow$16.48
80 49.69 40.26 42.49 $\uparrow$11.42 $\uparrow$8.52 $\uparrow$14.57
100 54.94 44.45 45.83 $\uparrow$16.67 $\uparrow$12.71 $\uparrow$17.91
Table 5: Few-shot learning ability. See scores of the validation set and more prompts in Section˜C.4.
6.6 Few-shot Learning Evaluation
@@
NatVLM undergoes instruction tuning with an extensive training dataset, DivSceneep. In this section, we design a few-shot experiment to confirm its ability to generalize with fewer data. While the original training data contains five episodes per house, we train the model with only 1, 2, 3, and 4 episodes, representing 20%, 40%, 60%, and 80% of the full data. The results are shown in Table˜5. For clarity, we also compare our models with the GPT-4o baseline. We can find that our agent has strong few-shot learning abilities. With only 20% percent training data, our agent can perform similarly to GPT-4o and generalize well on unseen houses. Meanwhile, the results demonstrate a gradual performance improvement as we incrementally increase the data volume, with performance plateauing at approximately 80% of the full dataset. We provide more analysis in Section˜C.4.
@@
$\diamond$ w/o ET 38.27 32.15 31.43 31.25 26.87 26.20
$\diamond$ w Diff-EQ 72.32 58.98 62.35 54.59 45.53 46.80
Table 6: Zero-shot transferring on iTHOR and ProcTHOR.
6.7 Zero-shot Transferring Evaluation
We further evaluate NatVLM on other house datasets: iTHOR (Weihs et al., 2021), ProcTHOR (Deitke et al., 2022), and HM3D Ramakrishnan et al. (2021). We directly use NatVLM tuned on DivSceneep to conduct zero-shot transferring evaluation. Since we do not tune hyperparameters on these datasets, we treat each dataset as a whole test set without any validation set. We show the results on iTHOR and ProcTHOR in Table˜6. We also report the performance of open-source VLMs and some ablated models as baselines. The results show that our agent surpasses all the baselines on both datasets, indicating that our framework has a strong ability to generalize in other environments. We include results on HM3D and setup details in Section˜C.5.
Figure 5: Error analysis. The agent needs to find a soda can in the lower right corner of a game room. The left image shows the predicted path, and the right image shows the ground truth in the grid map. The agent fails to find the soda can and wanders within a limited area.